A pyramid of height 15 ft is to be divided into three solids by passing
planes parallel to the base. Find the height of each solid if their volumes
are equal.
1
Expert's answer
2020-11-17T15:10:52-0500


The volume of whole this pyramide is:

"V=\frac{1}{3}Sh=\frac{1}{3}⋅S⋅15=5S"

According to the conditions of the problem

"V_1=V_2=V_3",

therefore

"V=V_1+V_2+V_3"

"V=3V_1"

"5S=3V_1"

"V_1=\frac{5S}{3}"

The volume of the top pyramide is:

"V_1=\frac{1}{3}S_1h_1", therefore

"\frac{5S}{3}=\frac{1}{3}S_1h_1"

"5S=S_1h_1"

"h_1=\frac{5S}{S_1}" - it is the height of the top pyramide

Then

"V_1+V_2=\frac{1}{3}S_2(h_1+h_2)"

"2V_1=\frac{1}{3}S_2(h_1+h_2)"

"2V_1=\frac{1}{3}S_2(\frac{5S}{S_1}+h_2)" "\/\u22c53"

"6V_1=S_2(\frac{5S}{S_1}+h_2)"

"\frac{5S}{S_1}+h_2=\frac{6V_1}{S_2}"

"h_2=\frac{6V_1}{S_2}-\frac{5S}{S_1}"

"h_2=\frac{6⋅\frac{5S}{3}}{S_2}-\frac{5S}{S_1}"

"h_2=\frac{10S}{S_2}-\frac{5S}{S_1}" - it is the height of the middle solide

"h_3=15-h_1-h_2=15-\frac{5S}{S_1}-(\frac{10S}{S_2}-\frac{5S}{S_1})="

"=15-\frac{5S}{S_1}-\frac{10S}{S_2}+\frac{5S}{S_1}=15-\frac{10S}{S_2}" - it is the height of the under solid


Solution: "\frac{5S}{S_1}" ; "\frac{10S}{S_2}-\frac{5S}{S_1}" ; "15-\frac{10S}{S_2}" .